# Portrait3D: 3D Head Generation from Single In-the-wild Portrait Image

### v1

@@ 4.2 Geometry Sculpting
During the geometry sculpting stage, we aim to obtain a fine-grained 3D geometry, that is consistent with the facial features in the input image under the reference view, and consistent across different viewpoints. We first convert the initial 3D mesh obtained from ID-aware initialization into a textured 3D mesh representation. In this process, we use deformable tetrahedral grid (DMTet) [40] as the 3D representation to facilitate high-resolution details and high-quality surface. Specifically, we use the normals rendered from 3D-GAN as the pseudo ground truth normals $\mathcal{N}_{\mathbf{c}}$. For the DMTet object to be optimized, we use Marching Tetrahedra to generate the normal map $\hat{\mathcal{N}_{\mathbf{c}}}$. In order to polish the overall surface from multiple viewpoints, we uniformly sample camera poses $\mathbf{c}$ distributed in the 3D-GAN space, and use $\mathcal{L}_{2}$ distance between normal maps as the loss function, defined as $\mathcal{L}_{\text{norm}}=\|\hat{\mathcal{N}_{\mathbf{c}}}-\mathcal{N}_{\mathbf{c}}\|_{2}$.
Geometry Refinement. During the geometry refinement stage, we use ID-aware Geometry Supervision (Sec. 4.1) to ensure the generation of fine geometry under the reference view, while supplementing the information with the prior of the diffusion model under other views, as shown in Eq.(4). Under the reference view, we calculate the reference loss from the head foreground mask, normal map, and depth map of the reference image, which is defined as:
@@ 4.3 ID-Consistent Texture Inpainting and Refinement
After obtaining a fine-grained 3D geometry in the geometry sculpting stage, during the texture generation stage, our goal is to rapidly generate realistic textures for the 3D head mesh. Previous texture generation methods [5, 44] often used SDS/VSD loss for multi-view optimization to obtain textures. However, the textures generated in this way may become over-saturated and inconsistent with the color tone of the reference image.
Therefore, we design an inpainting and refinement process to quickly produce high-quality UV texture maps without involving SDS/VSD optimization.
Progressive Texture Inpainting. With an uncolored 3D head surface mesh $M$ with vertices $V$ and triangular faces $F$ generated in the geometry generation stage, denoted as $M=(V,F)$, we initialize a UV texture map $\mathbf{T}_{0}$ corresponding to the mesh coordinates. Starting from the reference view $\mathbf{c}_{0}$, we first back-project the reference portrait image onto the 3D mesh. In order to maintain consistency with the reference image texture, we sample a camera trajectory $\{\mathbf{c}_{1},\mathbf{c}_{2},\cdots,\mathbf{c}_{n}\}$ (where $\mathbf{c}_{i}$ is the $i$-th viewpoint, and $n$ is the number of sampled viewpoints) to progressively generate texture information. The camera trajectory starts from the reference image viewpoint, gradually moves the camera pose to the back view of the head, and finally generates the texture for the top of the head that has not been generated yet.
During the inpainting process, at a specific viewpoint $\mathbf{c}_{k}$, given the texture map $\mathbf{T}_{k-1}$ that has been textured in $k-1$ previous viewpoints, we are able to render an RGB image $I_{k}$ with partially colored regions, as well as a mask $\mathbf{m}_{k-1}$ indicating the colored (visible) regions. To better inpaint the uncolored areas, we render a normal map $\boldsymbol{x}^{normal}_{k}$ from the 3D mesh as the geometric condition at the $\mathbf{c}_{k}$ viewpoint, and then use the normal-conditioned diffusion model to generate the texture for the uncolored areas, which can be formulated as:
@@
Texture Refinement. Although the UV texture map obtained through texture inpainting may exhibit inconsistent textures in some regions, it provides us with a good texture initialization. In the texture refinement stage, we aim to solve the texture inconsistency problem and further enhance the detail and realism of the generated texture.
We optimize the UV texture map by utilizing the ID-guided diffusion model (Sec. 4.1) and minimizing the distance between the rendered multi-view images and the refined images.
@@
$\hat{\boldsymbol{x}}_{0}=p^{\text{id}}_{0}(\boldsymbol{x}_{0}|\boldsymbol{x}_{t},y,f_{ID},\boldsymbol{x}^{normal})$ (11)
After obtaining the refined image $\hat{\boldsymbol{x}}_{0}$, we calculate the pixel-level MSE loss $\mathcal{L}_{\text{MSE}}$ between $\boldsymbol{x}_{0}$ and $\hat{\boldsymbol{x}}_{0}$ to further refine the texture map. Besides, we add a perceptual loss $\mathcal{L}_{\text{percep}}$ to enhance the texture detail information and maintain style similarity. We also calculate the MSE loss $\mathcal{L}_{\text{MSE}}^{\text{ref}}$ between $x_{0}$ from the reference view and the reference image $\boldsymbol{x}_{\mathrm{ref}}$. By combining these supervisions, we obtain the final loss $\mathcal{L}_{\text{refine}}$ for the texture refinement stage:
$\mathcal{L}_{\text{refine}}=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{percep}}+\mathcal{L}_{\text{MSE}}^{\text{ref}}.$ (12)
5 EXPERIMENT
Figure 5: Qualitative evaluation with image-to-3D methods. Our method surpasses previous approaches in terms of the facial detail in head geometry as well as the realism of the texture.
In the experimental section, we will start with the implementation details, followed by quantitative and qualitative comparisons between our method and others. Finally, we will analyze the role of each module in our method.
@@ 5.1 Implementation Details
Our Portrait3D is built upon the open-source project ThreeStudio [12]. In the geometry stage, we use SD1.5 [37] as the base diffusion model. During the texture generation stage, we employ Realistic Vision 4.0 as the base diffusion model, which is a version fine-tuned for real portraits on SD1.5. We set the resolution of DMTet to 512 to achieve better geometric details. Additionally, the size of the unfolded texture map from the mesh is set to 1024x1024. In the geometry generation stage, we perform 5,000 optimization iterations. During the texture generation stage, we first perform texture inpainting from 15 ordered viewpoints, followed by 400 steps of texture refinement. Our experiments are conducted on a machine with a single V100 GPU, with a batch size set to 1. For each 3D head, the total optimization time is approximately 1 hour.
Baselines. We conducted extensive comparisons with four other single-view 3D generation methods (Magic123[31], DreamCraft3D[41], Wonder3D[25]) and a single-view human head generation method (Panohead[1]). For Magic123, DreamCraft3D, and Wonder3D, we generated the 3D models strictly following the official open-source code and parameters. For Panohead, we followed the official guidance and obtained results using the GAN inversion method. In addition, we compared our method with the text-to-3D human head generation methods. For a fair comparison, we did not choose the non-open-source methods HeadSculpt[13] and HeadArtist[22]. Instead, we selected the state-of-the-art approach, HumanNorm[16], for it has excellent generation results, open-source code, and good reproducibility.
Datasets. We conducted experiments on a subset of facial images from the FFHQ [17] dataset, a high-quality in-the-wild facial dataset known for its rich diversity in age, ethnicity, and image backgrounds, as well as significant variations in facial attributes. This dataset presents a challenging task due to its unconstrained environment. From this dataset, We selected 106 portrait images with unoccluded faces to evaluate our model and baselines.
Evaluation metrics. We employ three commonly used metrics to assess the rendered images of the generated heads in terms of image quality and facial reconstruction. LPIPS [50] is utilized for fidelity measurement, while CLIP-I [47] score and ID [1] score are employed to gauge the multi-view consistency and identity preservation.
Figure 6: Qualitative evaluation with text-to-head method. Our method has a clear advantage in the realism of geometry and texture.
@@ 5.2 Qualitative Evaluation
Qualitative evaluation with image-to-3d methods. In Fig.5 we show the comparisons between our method and four other image-to-3D methods. These four images are selected from the in-the-wild face image dataset FFHQ and include faces of different ages and genders. For the geometry, the 3D generation methods based on optimization (Magic123 and Dreamcraft3D) can’t restore the stereoscopic shape of the human head. This is because they use view-dependent diffusion to guide the optimization process based on the 3D-SDS, and it does not generalize well to in-the-wild face images. Wonder3D can generate basic human head shapes but generally suffers from overly flat facial features in side views, and the facial surface lacks geometry detail. Besides, Panohead produces strange protrusions in the eye area. In contrast, our method, after introducing the identity information to the geometry sculpting stage, can restore fine geometry on the face while ensuring a high-quality overall head shape. For texture results, although Dreamcraft3D and Wonder3D can generate high-quality textures from the reference viewpoint, Dreamcraft3D, due to poor head geometry, is completely unable to generate images of the head from other angles. Wonder3D’s texture quality drops significantly on the side (i.e., the side rear view of the head in the first row of Wonder3D) and cannot produce high-resolution textures. Our method, after using ID-enhanced inpainting and refinement processes, can generate more photorealistic textures. In summary, our method can produce high-quality head geometry and realistic textures while fully restoring the facial information from the reference images.
Qualitative evaluation with text-to-head method. In Fig.6, we show a comparison of our method with the current state-of-the-art text-to-head method Humannorm. Since text-to-human head methods can only accept text as input and cannot fully describe in-the-wild face information, the generated head may not match the face in the reference image. Therefore, for a fairer comparison, we choose examples of celebrities so that the text-to-head method can accurately generate heads based on the celebrity names. For generated geometry, Humannorm tends to generate heads with exaggerated structures, whereas our method generates far more realistic geometries. In terms of texture, due to the use of SDS loss to optimize the texture, the heads generated by Humannorm often have over-saturated colors. In contrast, our method fully utilizes both the reference image information and the identity information of the person depicted to generate more realistic faces.
@@ 5.3 Quantitative Evaluation
For each generated head model, we render images from five consistent viewpoints: left-frontal, left, right-frontal, right, and rear. All five rendered images are included in computing the CLIP-I and LPIPS metrics alongside the reference image. In the calculation of the ID metric, we compute the average Adaface [18] cosine similarity score across the four viewpoints excluding the rear. Note that if facial landmarks cannot be detected in an image from a particular viewpoint, its score is set to 0. In Tab. 1, we present the quantitative comparisons with the previous SOTA methods. Our method significantly outperforms previous approaches in terms of the ID metric, indicating its effectiveness in capturing the facial features of individuals and demonstrating high identity consistency across various perspectives in the analysis of novel viewpoints.
User Study. To validate our model’s robustness and quality, we conducted a user study with 15 random examples. Participants ranked five methods’ outputs based on quality from two random viewpoints. Our model ranked highest on average across 480 responses from 32 participants, confirming its superior quality.
Table 1: Quantitative comparisons on head generation task.
@@
Magic123 0.5601 0.4447 0.0245 1.11
PanoHead 0.4343 0.6600 0.2737 4.09
Wonder3D 0.5095 0.6860 0.2729 3.13
DreamCraft3D 0.4941 0.6074 0.2526 1.95
Portrait3D (Ours) 0.4727 0.6605 0.4451 4.72
@@ 5.4 Ablation Study
Figure 7: Visual ablation result of our ID-aware Initialization and Guidance.
Figure 8: Visual ablation result of our ID-Consistent Texture Inpainting and Refinement. Our method effectively reduces the artifact in texture generation.
@@
Effectiveness of ID-aware Initialization and Guidance. We present the visual in Fig.7 result to show the effectiveness of our ID-aware Initialization and Guidance method. (a) By replacing the normal map obtained through ID-aware Geometry Supervision with the normal map extracted from DPT[35, 34], we can observe that without ID-aware Geometry Supervision, the generated geometry loses many facial details. (b) When removing the ID-aware Guidance Enhancement in the texture generation stage, we can see that the faces lose the constraints imposed by identity information and easily produce artifacts for large angles (e.g., side faces) without sufficient reference image information. (c) We compare our ID-aware Initialization results with the use of Flame as the initialization geometry. As Flame aligns only with the facial landmarks, the head portions other than the face in the in-the-wild images are difficult to align with the initialization geometry, leading to difficulty in optimizing the geometry.
Effectiveness of ID-Consistent Texture Inpainting and Refinement. As shown in Fig.8, without (a) Texture refinement stage, relying solely on inpainting can lead to discontinuous boundaries between each inpainting step, which is particularly noticeable at the intersection of the ears and hair. Additionally, if we do not adopt (b) Progressive Inpainting to complete the texture, but instead use a view-surrounding approach, incoherent texture colors will be generated (e.g. yellow part at the neck).
@@ 5.5 Application
ID-consistant Texture Stylization. With the powerful capability of text-to-image diffusion models, using stylized stable diffusion versions, we can modify the text prompt during the texture generation stage to achieve stylized 3D head generation. Due to the use of ID-aware guidance in the texture generation stage, the stylized heads generated can maintain the identity information of the character in the portrait image, as shown in Fig. 9.
@@ 6 CONCLUSION
We propose Portrait3D, a novel method to realize 3D head generation from a single in-the-wild portrait image. By adding identity knowledge to the initialization, guidance, and geometry supervision part of the optimization process, our method significantly improves ID consistency and geometric details. Moreover, we design a two-stage ID-consistent texture inpainting and refinement method, which leads to coherent and photo-realistic head texture. Portrait3D outperforms previous methods in both geometry and texture quality, demonstrating the ability to generate high-fidelity 3D heads. These advantages combined with stylized 3D head generation bring great help for personalized 3D head asset generation.
Limitations and future work. Although we can generate 3D human heads from portrait images, currently these 3D heads cannot be driven for tasks such as 3D talking avatars. In the future, we will attempt to develop drivable 3D human head generation methods.
Appendix
Appendix A Overview
In this supplementary material, we will show more content not mentioned in the main paper, including:

### v2

@@ 3.2 Geometry Sculpting
During the geometry sculpting stage, we aim to obtain a fine-grained 3D geometry, that is consistent with the input image under the reference view, and consistent across different viewpoints. We first convert the initial 3D mesh obtained from ID-aware initialization into DMTet (Shen et al. 2021) representation to facilitate high-resolution details and high-quality surface. Specifically, we use the normals rendered from 3D-GAN as the pseudo ground truth normals $\mathcal{N}_{\mathbf{c}}$. For the DMTet object to be optimized, we use Marching Tetrahedra to generate the normal map $\hat{\mathcal{N}_{\mathbf{c}}}$. To polish the overall surface from multiple viewpoints, we uniformly sample camera poses $\mathbf{c}$ distributed in the 3D-GAN space, and use $L_{2}$ distance between normal maps as the loss function, defined as $\mathcal{L}_{\text{norm}}=\|\hat{\mathcal{N}_{\mathbf{c}}}-\mathcal{N}_{\mathbf{c}}\|_{2}$.
Geometry Refinement. During the geometry refinement stage, we use ID-aware Geometry Supervision to ensure the generation of fine geometry under the reference view, while supplementing the information with the prior of the diffusion model under other views, as shown in Eq.(1). Under the reference view, we calculate the reference loss from the head foreground mask, normal map, and depth map of the reference image, which is defined as:
@@ 3.3 ID-Consistent Texture Inpainting and Refinement
During the texture generation stage, our goal is to rapidly generate realistic textures for the 3D head mesh. Previous textures generated from SDS/VSD loss (Chen et al. 2023; Wang et al. 2024) may become over-saturated and inconsistent with the color tone of the reference image. Therefore, we design an inpainting and refinement process to quickly produce high-quality UV texture maps.
Progressive Texture Inpainting. With an uncolored 3D head surface mesh $M$ with vertices $V$ and triangular faces $F$ generated in the geometry generation stage, denoted as $M=(V,F)$, we initialize a UV texture map $\mathbf{T}_{0}$ corresponding to the mesh coordinates. Starting from the reference view $\mathbf{c}_{0}$, we first back-project the reference portrait image onto the 3D mesh. To maintain consistency with the reference image texture, we sample a camera trajectory $\{\mathbf{c}_{1},\mathbf{c}_{2},\cdots,\mathbf{c}_{n}\}$ (where $\mathbf{c}_{i}$ is the $i$-th viewpoint, and $n$ is the number of sampled viewpoints) to progressively generate texture. The camera trajectory starts from the reference image viewpoint, gradually moves the camera pose to the back view of the head, and finally generates the texture for the top of the head that has not been generated yet.
During the inpainting process, at a specific viewpoint $\mathbf{c}_{k}$, given the texture map $\mathbf{T}_{k-1}$ that has been textured in $k-1$ previous viewpoints, we can render an RGB image $I_{k}$ with partially colored regions, as well as a mask $\mathbf{m}_{k-1}$ indicating the colored (visible) regions. To better inpaint the uncolored areas, we render the normal map $\boldsymbol{x}^{normal}_{k}$ from the 3D mesh as the geometric condition at the $\mathbf{c}_{k}$ viewpoint, and then use the normal-conditioned diffusion model to generate the texture for the uncolored areas, which is formulated as:
@@
Figure 5: Qualitative evaluation with image-to-3D methods. Our method surpasses previous approaches in terms of the facial detail in head geometry as well as the realism of the texture. The zoom-in result of eye/mouth region and facial geometry shows that our method has clearer structural details and more reasonable geometric structure.
Texture Refinement. In the texture refinement stage, we aim to solve the texture inconsistency problem that appears in texture inpainting and further enhance the detail and realism of the generated texture.
We optimize the UV texture map by utilizing the ID-guided diffusion model and minimizing the distance between the rendered images and the refined images.
@@
$\hat{\boldsymbol{x}}_{0}=p^{\text{id}}_{0}(\boldsymbol{x}_{0}|\boldsymbol{x}_{t},y,f_{ID},\boldsymbol{x}^{normal}).$ (8)
After obtaining the refined image $\hat{\boldsymbol{x}}_{0}$, we calculate the pixel-level MSE loss $\mathcal{L}_{\text{MSE}}$ between $\boldsymbol{x}_{0}$ and $\hat{\boldsymbol{x}}_{0}$ to further refine the texture map. Besides, we add a perceptual loss $\mathcal{L}_{\text{percep}}$ to enhance the texture detail information and maintain style similarity. We also calculate the MSE loss $\mathcal{L}_{\text{MSE}}^{\text{ref}}$ and perceptual loss $\mathcal{L}_{\text{percep}}^{\text{ref}}$ between $x_{0}$ from the reference view and the reference image $\boldsymbol{x}_{\mathrm{ref}}$. By combining these supervisions, we obtain the final loss $\mathcal{L}_{\text{refine}}$ for the texture refinement stage:
$\mathcal{L}_{\text{refine}}=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{percep}}+\mathcal{L}_{\text{MSE}}^{\text{ref}}+\mathcal{L}_{\text{percep}}^{\text{ref}}.$ (9)
4 Experiments
In the experimental section, we start with the implementation details, followed by quantitative and qualitative comparisons between our method and others. Finally, we analyze the role of each module in our method.
@@ 4.1 Implementation Details
Our ID-Sculpt is built upon the open-source project ThreeStudio (Guo et al. 2023). In the geometry stage, we use SD1.5 (Rombach et al. 2022) as the base diffusion model. During the texture generation stage, we employ Realistic Vision 4.0 as the base diffusion model. In the geometry generation stage, we perform 5,000 optimization iterations. During the texture generation stage, we first perform texture inpainting from 15 ordered viewpoints, followed by 400 steps of texture refinement. Our experiments are conducted on a single V100 GPU, with a batch size set to 1. For each 3D head, the total optimization time is approximately 40 minutes.
Baselines. We conducted extensive comparisons with four other single-view 3D generation methods (Magic123 (Qian et al. 2023), DreamCraft3D(Sun et al. 2023), Wonder3D (Long et al. 2023)) and a single-view human head generation method (Panohead (An et al. 2023)). For Magic123, DreamCraft3D, and Wonder3D, we generated the 3D models strictly following the official open-source code and parameters. For Panohead, we followed the official guidance and obtained results using the GAN inversion method. In addition, we compared our method with the state-of-the-art text-to-3D human head generation methods HumanNorm (Huang et al. 2023), for it has excellent generation results and good reproducibility.
Datasets. We conducted experiments on a subset of facial images from the FFHQ (Karras, Laine, and Aila 2019) dataset, a high-quality in-the-wild facial dataset known for its rich diversity in age, ethnicity, and image backgrounds, as well as significant variations in facial attributes. This dataset presents a challenging task due to its unconstrained environment. From this dataset, We selected 106 portrait images with unoccluded faces to evaluate our model and baselines.
Evaluation metrics. We employ LPIPS (Zhang et al. 2018) for fidelity measurement, CLIP-I (Ye et al. 2023) score and ID (An et al. 2023) score to measure the multi-view consistency and identity preservation.
Figure 6: Qualitative evaluation with text-to-head method. Our method has a clear advantage in the realism.
@@ 4.2 Qualitative Evaluation
Qualitative evaluation with image-to-3D methods. In Fig. 5 we show the comparisons between our method and four other image-to-3D methods. These four images are selected from the in-the-wild face image dataset FFHQ and include faces of different ages and genders. For the geometry, the 3D generation methods based on optimization (Magic123 and Dreamcraft3D) can’t restore the stereoscopic shape of the human head. Wonder3D can generate basic human head shapes but generally suffers from overly flat facial features in side views, and the facial surface lacks geometry detail. Besides, Panohead produces strange protrusions in the eye area. In contrast, our method, after introducing the identity information to the geometry sculpting stage, can restore fine geometry on the face while ensuring a high-quality overall head shape. For texture results, Dreamcraft3D and Wonder3D cannot generate images of the head from other angles. Wonder3D’s texture quality drops significantly on the side (i.e., the side rear view of the head in the first row of Wonder3D) and cannot produce high-resolution textures. Panohead struggles to generate high-fidelity structural details in areas like the eye and mouth region. After using ID-enhanced inpainting and refinement processes, our method can generate more photorealistic textures. In summary, our method can produce high-quality head geometry and realistic textures while fully restoring the facial information from the reference images.
Qualitative evaluation with text-to-head method. In Fig. 6, we show a comparison of our method with the current SOTA text-to-head method HumanNorm. Since text description cannot fully describe in-the-wild face information, we choose examples of celebrities so that the text-to-head method can accurately generate heads based on the celebrity names. For generated geometry, HumanNorm tends to generate heads with exaggerated structures, whereas our method generates far more realistic geometries. In terms of texture, the heads generated by HumanNorm have over-saturated colors. In contrast, our method fully utilizes both the reference image information and the identity information of the person depicted to generate more realistic faces.
@@ 4.3 Quantitative Evaluation
Table 1: Quantitative comparisons on head generation task. We compute LPIPS under the reference view, ClIP-I and ID score under novel views.
@@
Magic123 0.367 0.4447 0.0245 1.11
PanoHead 0.045 0.6600 0.2737 4.09
Wonder3D 0.135 0.6860 0.2729 3.13
DreamCraft3D 0.068 0.6074 0.2526 1.95
ID-Sculpt (Ours) 0.039 0.6605 0.4451 4.72
For each generated head model, we render images from four consistent viewpoints: left-frontal, left, right-frontal, and right. In Tab. 1, we present the quantitative comparisons with the previous SOTA methods. Our method outperforms previous approaches in terms of the ID metric and LPIPS score, indicating its effectiveness in capturing facial features and demonstrating high identity consistency across various perspectives in the analysis of novel viewpoints.
User Study. To validate our model’s robustness and quality, we conducted a user study with 15 random examples. Participants ranked five methods’ outputs based on quality from two random viewpoints. Our model ranked highest on average across 480 responses from 32 participants.
@@ 4.4 Ablation Study
Figure 7: Visual ablation results of our ID-aware Initialization and Guidance.
Figure 8: Visual ablation results of our ID-Consistent Texture Inpainting and Refinement. Our method effectively reduces the artifacts in texture generation.
@@
Effectiveness of ID-aware Initialization and Guidance. We present the visual result in Fig.7 result to show the effectiveness of our ID-aware Initialization and Guidance method. (a) By replacing the normal map obtained through ID-aware Geometry Supervision with the normal map extracted from DPT(Ranftl, Bochkovskiy, and Koltun 2021; Ranftl et al. 2020), we can observe that without ID-aware Geometry Supervision, the generated geometry loses many facial details. (b) When removing the ID-aware Guidance Enhancement in the texture generation stage, we can see that the faces lose the constraints imposed by identity information and easily produce artifacts for large angles (e.g., side faces) without sufficient reference image information. (c) We compare our ID-aware Initialization results by using Flame as the initialization geometry. As Flame aligns only with the facial landmarks, the head portions other than the face in the images are difficult to align with the initialization geometry, leading to difficulty in geometry optimization.
Effectiveness of ID-Consistent Texture Inpainting and Refinement. As shown in Fig.8, without (a) Texture refinement=, relying solely on inpainting can lead to discontinuous boundaries between each inpainting step, which is particularly noticeable at the intersection of the ears and hair. Without (b) Progressive Inpainting to complete the texture and instead use a view-surrounding approach, incoherent texture colors will be generated (e.g. yellow part at the neck).
@@ 4.5 Application
ID-consistant Texture Stylization. With the capability of stylized text-to-image diffusion models, we achieve stylized 3D head generation. Due to the use of ID-aware guidance in the texture generation stage, the generated stylized heads can maintain the identity information of the character in the portrait image, as shown in Fig 9.
@@ 5 Conclusion
We propose ID-Sculpt, a novel method to generate a 3D head from a single in-the-wild portrait image. Our method significantly improves ID consistency and geometric details by adding identity knowledge to the initialization, guidance, and geometry supervision part of the optimization process. Moreover, we design a two-stage ID-consistent texture inpainting and refinement method, which leads to coherent and photo-realistic head texture. ID-Sculpt outperforms previous methods in both geometry and texture quality, demonstrating the ability to generate high-fidelity 3D heads. These advantages combined with stylized 3D head generation bring great help for personalized 3D head asset generation.
Limitations and future work. Although we can generate high-quality 3D heads from a portrait image, the optimization-based approach takes the whole generation a long time (40 minutes). In the future, we will work on accelerating the speed of 3D head generation.
Acknowledgments
This work was supported by National Natural Science Foundation of China (No. 72192821, 62302297, 62472282), Shanghai Sailing Program (22YF1420300), Young Elite Scientists Sponsorship Program by CAST (2022QNRC001), the Fundamental Research Funds for the Central Universities (project number: YG2023QNA35).
Appendix A Appendix Overview
In the appendix, we will show more content not mentioned in the main paper, including: